OpenStudy (anonymous):

what is each expression written using each base only once 8^8 . 8^4 A.512^12 B.8^32 C.24^12 D.8^12

jimthompson5910 (jim_thompson5910):

hint: \[\Large a^b*a^c = a^{b+c}\]

jimthompson5910 (jim_thompson5910):

For example, \[\Large 2^3*2^5 = 2^{3+5} = 2^8\]

OpenStudy (anonymous):

so it would 8^12?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

could you help me with another questions

jimthompson5910 (jim_thompson5910):

sure I can help with one more

OpenStudy (anonymous):

ok this one is in a drawing

OpenStudy (anonymous):